# What Foundation Models can Bring for Robot Learning in Manipulation : A Survey

### v2

@@ 1 Introduction
Foundation models are primarily pretrained on vast internet-scale datasets, enabling them to be fine-tuned for diverse tasks. Their significant advancements in vision and language processing contribute to mitigating the aforementioned challenges. Based on the different input modalities and functionalities of the models, we categorize foundation models into the following six types.
@@
    Vision Foundation Models (VFMs) like SAM (Kirillov et al. (2023)) demonstrate strong segmentation capability for open-set objects.
@@
    Visual Content Generation Models (VGMs) like DALL-E (Ramesh et al. (2021)), Zero-1-to-3 (Liu et al. (2023c)), demonstrate the capability to generate 2D images or 3D meshes through text or images.
@@
    Vision Language Models (VLMs) like GPT-4V (Achiam et al. (2023)), CLIP (Radford et al. (2021)) showcase robust comprehension of both vision and language, such as open-set image classification and visual question answering.
@@
    Large Multimodal Models (LMMs) expand their scope beyond vision and language to create novel categories of foundation models incorporating additional modalities, such as ULIP (Xue et al. (2023a)) aligns point cloud representation to the pre-aligned image-text feature space.
@@
    Robot Foundation Models (RFMs), like RT-X (Padalkar et al. (2023a)). Internet-scale dataset, such as images and text data, are suitable for pre-training visual and language models, but lack task level manipulation data. Therefore, researchers aim to train end-to-end RFMs by collecting task-level manipulation datasets to enable observations-to-action mapping.
@@
    RFMs serve as an end-to-end policy model, directly outputting actions based on input blue observations.
@@
These findings underscore the potential of embedding foundation models into manipulation tasks as a viable path toward achieving general manipulation capability. However, we do not believe that a single foundation model alone can achieve general manipulation capability. Although RFMs currently represent a single-model end-to-end training approach, ensuring safety and stability, particularly in achieving an over 99% success rate in manipulation tasks, remains a challenge. Achieving over a 99% success rate in manipulation tasks is crucial, as human manipulation success rates are around 99%. Without this level of accuracy, robots can’t replace humans. Therefore, drawing inspiration from the development of autonomous driving systems (Hu et al. (2023c)), achieving general manipulation capability necessitates an overarching framework that encompasses multiple functional modules, with different foundation models assuming distinct roles in facilitating general manipulation capability.
The ultimate general manipulation framework should be able to interact with human or other agent and to manipulate arbitrary objects in open-world scenarios and achieve diverse manipulation tasks. However, the interaction between robot and human involves not only recognizing intentions but also learning new skills or improving old skills from human experts in the external world. Open-world scenarios may be static or dynamic. Objects can be either rigid or deformable. Task objectives can vary from short-term to long-term. Furthermore, tasks may necessitate different degrees of precision with respect to contact points and applied forces/torques. We designate the restriction of the robot’s learning capability to improving old skills and to manipulating rigid objects in static scenes in order to achieve short-horizon task objectives with low precision requirements for contact points and forces/torques as Level 0 (L0). Drawing from Kroemer et al. (2021), we propose a comprehensive framework for general manipulation task at L0. Hence, we aim to use this survey not only to enlighten scholars on the issues that foundation models can address in robot learning for manipulation but also to stimulate their exploration of the general manipulation framework and the role various foundation models can play in the general manipulation framework.
@@ 2 Framework of Robot Learning for General Manipulation
Early stage, robot manipulation is defined as learning a policy $\Pi$ through deep learning, reinforcement learning, or imitation learning, etc. This policy controls the robot’s joint movements and executes tasks based on observations of the environment and the robot’s state $S$, mapping to actions $\alpha$. such as Rlafford (Geng et al. (2023)) and Graspnet (Fang et al. (2020)) take point cloud as input and output the target pose. This process is represented by the Skill Execution module, as shown in Fig. 2.
@@
Recently, researchers have realized that training policies require real-world interaction between the robot and environments, which inevitably increases the probability of unforeseen hazardous situations. Therefore, researchers aim to train the environment’s transition model $T$. Once the model is fitted, robot can generate samples based on it, significantly reducing the frequency of direct interaction between the robot and environments (Liu et al. (2024)). This process is represented as the Transition module $T$ in Fig. 2.
@@
    Transition. The Transition module $T$ takes an action generated by the Policy module $P$ as input. It outputs the next state after executing this action, thus helping to reduce the interaction between the robot and the real environment. Zhang and Soh (2023) uses LLM as a human model, predicting the changes in human state after robot actions, and selecting the most suitable robot actions for interaction with humans. In various algorithms, the transition environment model is not a mandatory module. Therefore, in Fig. 2, we enclose this module with a dashed box.
@@ 3 Human/Agent Interaction
In addressing instruction ambiguity and text generation and comprehension challenges, SeeAsk (Mo et al. (2023)) utilizes CLIP’s perceptual module to identify objects in the scene and employs a fixed questioning template to organize language to ask about which object will be manipulated. Although the use of CLIP enhances the generalization ability for object recognition, it can’t generate text for asking questions and to comprehend answers from the outside world and SeeAsk (Mo et al. (2023)) focuses solely on addressing ambiguities concerning object color and spatial relationship due to a fixed questioning template. KNOWNO (Ren et al. (2023a)) utilizes LLM to score the next action to be taken. If the score difference between the top two actions is less than a threshold, it’s considered ambiguity, prompting a confirmation for the final action. This approach improves efficiency and autonomy. Matcha (Zhao et al. (2023b)) not only employs vision but also utilizes haptic and sound senses to perceive object properties, such as material. When encountering ambiguity in object attribute recognition, it leverages LLM to generate inquiry content. CoELA (Zhang et al. (2023b)) utilizes LLM as both a communication module and a planning module to enhance interaction text generation and comprehension, as well as task scheduling, with collaborative agent. LLM-GROP (Ding et al. (2023)) utilizes LLM to extract latent commonsense knowledge embedded within task instruction. For example, a task instruction might be ”set dinner table with plate and fork,” while the latent commonsense knowledge could be ”fork is on the left of a bread plate.”
@@ 4 Pre- and Post-conditions Detection
In pre- and post-conditions detection, it is necessary to identify the initial and termination conditions. In pre-conditions detection, recognize objects and observe the affordances of objects. Object recognition will be addressed in the section 6. In post-conditions detection, identify whether a task has been successfully executed and provide reasons for task failure after skill execution. Currently, there are few papers focusing on identifying termination conditions. Cui et al. (2022) utilizes CLIP to compare the target’s text or image with the termination environment to determine the success of task execution. Few articles are found in this study that address the output of task failure reasons after skill execution. RobotGPT (Jin et al. (2024)) analysis task failure utilizes the positions of manipulated objects after execution, but task failure should be determined during execution. Therefore, this section focuses on literature discussing foundation models in object affordance.
4.1 Affordance
@@
The methods mentioned above utilize foundation models to assist other learning methods in improving affordance maps or grasp poses. There are also direct approaches using foundation models to generate affordance maps and grasp poses. PartSLIP (Liu et al. (2023b)) converts 3D point clouds into 2D rendering images and inputs multi-view 2D images and textual descriptions of object parts into GLIP for object parts detection, ultimately fusing 2D bounding boxes into 3D segmentation to generate affordance maps. LAN-grasp (Mirjalili et al. (2023)) inputs human instruction into LLM, utilizing its prior knowledge to output the shape of part to be grasped. These shapes, along with the object’s 2D image, are then inputted into VLM to detect the bounding box for the grasping part. Finally, the bounding box and the point cloud from object 3D reconstruction are inputted into the grasp planner to generate grasp poses.
@@
Following Fig. 4, current research explores LLMs’ zero/few-shot capability in providing object part-level prior knowledge via text, aiding in generating affordance maps or grasp poses. Direct use of foundation models for affordance is possible but still shows instability, with unclear performance boundaries.
@@
Figure 4: Foundation Models for Affordance. The main approaches of task-oriented grasp are supervised learning and reinforcement learning. Both methods utilize LLM to generate object part-level description and desired affordance description in task instruction, then fuse tokens and features into the original network through language encoder and image encoder to output task-oriented grasp pose (Tang et al. (2023); Ren et al. (2023b)). In reinforcement learning, it is possible to choose between a LLM language encoder with a custom-designed image encoder, or a VLM language encoder with a VLM image encoder. When selecting the LLM language encoder with a custom image encoder, the LLM language encoder should be frozen, and the custom image encoder should be trained. When using the VLM language encoder with the VLM image encoder, both encoders are typically frozen. Direct using foundation method utilizes LLM to generate object part-level description and desired affordance description according to task instruction. VLM marks out the part of the object to grasp in the image based on the description (Liu et al. (2023b))
@@ 5 Hierarchy of skills
In the last two decades, research on skill hierarchy has led to powerful domain-independent planners and various real-world applications (Silver et al. (2022)). Models for skill hierarchy can be trained using text or videos, similar to how humans learn assembly procedures from instructional manuals or tutorial videos, such as VLaMP (Patel et al. (2023)), which trains models to understand human video operations. Currently, the text-based approach dominates the skill hierarchy domain. PDDL, a Lisp-like language (Silver et al. (2022)), is commonly used to address skill hierarchy issues. However, as LLMs excel in natural language tasks and PDDL is not a natural language, researchers are exploring how LLMs can be employed by robot for skill hierarchy tasks (Vemprala et al. (2023); Jansen (2020); Driess et al. (2023)). Additionally, various benchmarks such as PlanBench (Valmeekam et al. (2023)) and CALVIN (Mees et al. (2022)) can assess the planning and reasoning capability of LLMs.
@@
Regarding user preferences, TidyBot (Wu et al. (2023b)) trains LLM by collecting users’ preference data, enabling the trained LLM to choose behaviors that better align with user preferences. As for safety regulations, Yang et al. (2023d) incorporate ISO 61508, a global standard for safely deploying robots in industrial factory settings, into the constraints of the action sequence generation. As for the skill library, BOSS (Zhang et al. (2023c)) suggests using LLMs’ rich knowledge to guide skills chaining in the skill library, aiming to create new skills through combinations. RoboGen (Wang et al. (2023b)) employs generative models to create new skill task scenarios, then utilizes either reinforcement learning or gradient optimization methods to automatically learn new skills based on the reward function generated by the LLM. As for action execution feedback, REACT (Yao et al. (2022b)), COWP (Ding et al. (2022)), LLM-Planner (Song et al. (2023)), CoPAL (Joublin et al. (2023)) provide feedback on robot action execution to LLMs. This allows LLMs to adjust action sequences based on execution status, creating a closed-loop process for generating action sequences.
As for the limitation of scene grounding, LLMs need to inquire about the scene representation to determine the availability, relationship and location of objects. NLMap (Chen et al. (2023a)) proposes an open-vocabulary, queryable semantic representation map built on ViLD and CLIP. This map outputs the pose of related objects based on task instruction, which are then handed over to the LLM for planning. Text2Motion (Lin et al. (2023)) incorporates a geometric value function on top of the value function, enabling the robot to select actions that adhere to geometric constraints based on scene descriptions. Xu et al. (2023b) explore the possibility of teaching robots to creatively utilize tools within scenarios, which involve implicit physical limitations and require long-term planning. VILA (Hu et al. (2023a)) seamlessly incorporates perceptual data into ChatGPT-4V for its reasoning and planning processes, facilitating a deep comprehension of common sense knowledge within the visual domain, encompassing spatial arrangements and object characteristics. PHYSOBJECTS (Gao et al. (2023)) fine-tunes a VLM to enhance its understanding of physical object attributes, such as material. This integration of a physically informed VLM into an interactive framework with a LLM enhances task planning performance in tasks incorporating instruction related to physical object attributes.
@@
6 State Perception
Compared to passive perception, active perception adjusts the perspective to the areas of interest (Kroemer et al. (2021)). Then, modeling manipulation tasks and generalizing manipulation skills necessitate representations of both the robot’s environment and the manipulated objects. These representations form the foundation for skill hierarchies, pre- and post-condition detection, skill learning, and transition model learning. The Vision Transformers (ViTs) and similar attention-based neural networks have recently achieved state-of-the-art performance on numerous computer vision benchmarks (Han et al. (2022); Khan et al. (2022); Zhai et al. (2022)) and the scaling of ViTs has driven breakthrough capability for vision models (Dehghani et al. (2023)).
Pre-trained visual representations have high transfer ability to policy learning (Xiao et al. (2022b); Yang et al. (2023c)), but visual representation involves not just recognizing spatial features but also understanding semantic features. Masked autoencoding methods prioritize low-level spatial aspects, sacrificing high-level semantics, whereas contrastive learning methods focus on the inverse (Karamcheti et al. (2023)). The fusion of masked autoencoder and contrastive learning is employed in both Voltron (Karamcheti et al. (2023)) and iBOT (Zhou et al. (2021)). The loss function achieves a balanced trade-off between these two aspects. To compare different pre-trained visual representations, benchmarks are established by CORTEXBENCH (Majumdar et al. (2023)) and EmbCLIP (Khandelwal et al. (2022)) to assess which model could provide a better ”artificial visual cortex” for manipulation tasks. However, the models included in these benchmarks are still not comprehensive enough. At the same time, it is important to consider the potential of the diffusion model (Ze et al. (2023)). The aforementioned pre-trained visual representations mainly involve the extraction of features from 2D images. The experience of learning representations on 2D images can also be extended to other modalities. For the object point cloud modality, ULIP (Xue et al. (2023a)) and ULIP2 (Xue et al. (2023b)) employ contrastive learning to align features between point clouds and text-images. Point-BERT (Yu et al. (2022)) uses the masked autoencoding method to learn point cloud features by reconstructing point clouds. In the haptic modality, MOSAIC (Tatiya et al. (2023)) utilizes contrastive learning to train the haptic encoder.
As for segmentation, SAM (Kirillov et al. (2023)) develops a transformer-based architecture and creates the largest segmentation dataset, with over 1 billion masks from 11 million images. The model is adaptable and enables zero-shot transfer to new tasks and image distributions. Fast-SAM (Zhao et al. (2023a)) and Faster-SAM (Zhang et al. (2023a)) aim to improve the training and inference speed of the network by enhancing its network structure. TAM (Yang et al. (2023a)) merges SAM (Yang et al. (2023a)) and XMem (Cheng and Schwing (2022)) for high-performance interactive tracking and segmentation in videos.
As for detection, traditional detection models are usually confined to a narrow range of semantic categories because of the cost and time involved in gathering localized training data within extensive or open-label domains. However, advancements in language encoders and contrastive image-text training enable open-set detection. Researchers integrate language into a closed-set detector to generalize open-set concepts, detecting various classes through language generalization despite being trained solely on existing bounding box annotations, such as OWL-ViT (Minderer et al. (2022)), Grounding-DINO (Liu et al. (2023d)), OVD (Zareian et al. (2021)), ViLD (Gu et al. (2021)), DetCLIP (Yao et al. (2022a)).
The method for open-set detection on 2D images can be extended to the research direction of open-set detection on 3D point clouds. PointCLIP (Zhang et al. (2022)) utilizes pre-trained CLIP to extract multi-view depth image features of point cloud, then compares the extracted features with language features to identify the point cloud category.

### v3

@@ 1 Introduction
Foundation models are primarily pretrained on vast internet-scale datasets, enabling them to be fine-tuned for diverse tasks. Their significant advancements in vision and language processing contribute to mitigating the aforementioned challenges. Based on Firoozi et al. (2023) and considering the different input modalities and functionalities of the models, we categorize foundation models into the following six types.
@@
    Visual Foundation Models (VFMs) like SAM (Kirillov et al. (2023)) demonstrate strong segmentation capability for open-set objects.
@@
    Visual Generative Models (VGMs) like DALL-E (Ramesh et al. (2021)), Zero-1-to-3 (Liu et al. (2023d)) and Sora (Brooks et al. (2024)), demonstrate the capability to generate 2D images, videos or 3D meshes through text or images.
@@
    Visual-Language Models (VLMs) like GPT-4V (Achiam et al. (2023)), CLIP (Radford et al. (2021)) showcase robust comprehension of both vision and language, such as open-set image classification and visual question answering.
@@
    Large Multimodal Models (LMMs) expand their scope beyond vision and language to create novel categories of foundation models incorporating additional modalities, such as ULIP (Xue et al. (2023a)) aligns point cloud representation to the pre-aligned image-text feature space. VLMs are a type of LMMs (Firoozi et al. (2023)). However, due to the current literature focusing more on VLMs, this paper will treat VLMs as a separate category. To avoid confusion, LMMs in this paper refer to those that include images, language, and more modalities.
@@
    Robotic-specific Foundation Models (RFMs), like RT-X (Padalkar et al. (2023a)). Internet-scale dataset, such as images and text data, are suitable for pre-training visual and language models, but lack task level manipulation data. Therefore, researchers aim to train end-to-end RFMs by collecting task-level manipulation datasets to enable observations-to-action mapping.
@@
    RFMs serve as an end-to-end policy model, directly outputting actions based on input observations.
@@
These findings underscore the potential of embedding foundation models into manipulation tasks as a viable path toward achieving general manipulation capability. However, we do not believe that a single foundation model alone can achieve general manipulation capability. Although RFMs currently represent a single-model end-to-end training approach, ensuring safety and stability, particularly in achieving an over 99% success rate in manipulation tasks, remains a challenge. Achieving over a 99% success rate in manipulation tasks is crucial, as human manipulation success rates are around 99%. Without this level of accuracy, robots can’t replace humans (Kumar (2023)). Therefore, drawing inspiration from the development of autonomous driving systems (Hu et al. (2023c)), achieving general manipulation capability necessitates an overarching framework that encompasses multiple functional modules, with different foundation models assuming distinct roles in facilitating general manipulation capability.
The ultimate general manipulation framework should be able to interact with human or other agent and to manipulate arbitrary objects in open-world scenarios and achieve diverse manipulation tasks (McCarthy et al. (2024)). Drawing from Kroemer et al. (2021) and this general manipulation definition, we propose a comprehensive framework for general manipulation. However, the interaction between robot and human involves not only recognizing intentions but also learning new skills or improving old skills from human experts in the external world. Open-world scenarios may be static or dynamic. Objects can be either rigid or deformable. Task objectives can vary from short-term to long-term. Furthermore, tasks may necessitate different degrees of precision with respect to contact points and applied forces/torques. Although there are many challenges, achieving general manipulation can be accomplished through multiple stages. We designate the restriction of the robot’s learning capability to improving old skills and to manipulating rigid objects in static scenes in order to achieve short-horizon task objectives with low precision requirements for contact points and forces/torques as Level 0 (L0). At the same time, we believe that improving the algorithm performance of different modules in the framework can support the transition from the L0 stage to the final general manipulation. Hence, we aim to use this survey not only to enlighten scholars on the issues that foundation models can address in robot learning for manipulation but also to stimulate their exploration of the general manipulation framework and the role various foundation models can play in the general manipulation framework.
@@ 2 Framework of Robot Learning for General Manipulation
Early stage, robot manipulation is defined as learning a policy $\Pi$ through deep learning, reinforcement learning, or imitation learning, etc. This policy controls the robot’s joint movements and executes tasks based on observations of the environment and the robot’s state $S$, mapping to actions $\alpha$. such as Rlafford (Geng et al. (2023b)) and Graspnet (Fang et al. (2020b)) take point cloud as input and output the target pose. This process is represented by the Skill Execution module, as shown in Fig. 2.
@@
Recently, researchers have realized that training policies require real-world interaction between the robot and environments, which inevitably increases the probability of unforeseen hazardous situations. Therefore, researchers aim to train the environment’s transition model $T$. Once the model is fitted, robot can generate samples based on it, significantly reducing the frequency of direct interaction between the robot and environments (Liu et al. (2024e)). This process is represented as the Transition module $T$ in Fig. 2.
@@
    Transition. The Transition module $T$ takes an action generated by the Policy module $P$ as input. It outputs the next state after executing this action, thus helping to reduce the interaction between the robot and the real environment. UniSim (Yang et al. (2023b)) introduces the action-in-video-out framework as an observation prediction model. It takes the current action as input and produces the subsequent observation as output.
@@ 3 Human/Agent Interaction
In addressing instruction ambiguity and text generation and comprehension challenges, SeeAsk (Mo et al. (2023)) utilizes CLIP’s perceptual module to identify objects in the scene and employs a fixed questioning template to organize language to ask about which object will be manipulated. Although the use of CLIP enhances the generalization ability for object recognition, it can’t generate text for asking questions and to comprehend answers from the outside world and SeeAsk (Mo et al. (2023)) focuses solely on addressing ambiguities concerning object color and spatial relationship due to a fixed questioning template. KNOWNO (Ren et al. (2023a)) utilizes LLM to score the next action to be taken. If the score difference between the top two actions is less than a threshold, it’s considered ambiguity, prompting a confirmation for the final action. This approach improves efficiency and autonomy. Matcha (Zhao et al. (2023c)) not only employs vision but also utilizes haptic and sound senses to perceive object properties, such as material. When encountering ambiguity in object attribute recognition, it leverages LLM to generate inquiry content. CoELA (Zhang et al. (2023b)) utilizes LLM as both a communication module and a planning module to enhance interaction text generation and comprehension, as well as task scheduling, with collaborative agent. LLM-GROP (Ding et al. (2023)) utilizes LLM to extract latent commonsense knowledge embedded within task instruction. For example, a task instruction might be ”set dinner table with plate and fork,” while the latent commonsense knowledge could be ”fork is on the left of a bread plate.”
@@ 4 Pre- and Post-conditions Detection
In pre- and post-conditions detection, it is necessary to identify the initial and termination conditions. In pre-conditions detection, recognize objects and observe the affordances of objects. In post-conditions detection, identify whether a task has been successfully executed and provide reasons for task failure after skill execution. Currently, there are few papers focusing on identifying termination conditions. Cui et al. (2022) utilizes CLIP to compare the target’s text or image with the termination environment to determine the success of task execution. Few articles are found in this study that address the output of task failure reasons after skill execution. RobotGPT (Jin et al. (2024)) analysis task failure utilizes the positions of manipulated objects after execution, but task failure should be determined during execution. Therefore, this section focuses on literature discussing foundation models in pre-conditions detection including object affordance and object recognition.
4.1 Object Affordance
@@
The methods mentioned above utilize foundation models to assist other learning methods in improving affordance maps or grasp poses. There are also direct approaches using foundation models to generate affordance maps and grasp poses. PartSLIP (Liu et al. (2023c)) converts 3D point clouds into 2D rendering images and inputs multi-view 2D images and textual descriptions of object parts into GLIP for object parts detection, ultimately fusing 2D bounding boxes into 3D segmentation to generate affordance maps. LAN-grasp (Mirjalili et al. (2023)) inputs human instruction into LLM, utilizing its prior knowledge to output the shape of part to be grasped. These shapes, along with the object’s 2D image, are then inputted into VLM to detect the bounding box for the grasping part. Finally, the bounding box and the point cloud from object 3D reconstruction are inputted into the grasp planner to generate grasp poses.
4.2 Object Recognition
Object recognition can be categorized into two types: passive perception and active perception. Compared to passive perception, active perception adjusts the perspective to the areas of interest (Kroemer et al. (2021)). Then, modeling manipulation tasks and generalizing manipulation skills necessitate representations of both the robot’s environment and the manipulated objects. These representations form the foundation for skill hierarchies, pre- and post-condition detection, skill learning, and transition model learning. The Vision Transformers (ViTs) and similar attention-based neural networks have recently achieved state-of-the-art performance on numerous computer vision benchmarks (Han et al. (2022); Khan et al. (2022); Zhai et al. (2022)) and the scaling of ViTs has driven breakthrough capability for vision models (Dehghani et al. (2023)).
Pre-trained visual representations have high transfer ability to policy learning (Xiao et al. (2022b); Yang et al. (2023c)), but visual representation involves not just recognizing spatial features but also understanding semantic features. Masked autoencoding methods prioritize low-level spatial aspects, sacrificing high-level semantics, whereas contrastive learning methods focus on the inverse (Karamcheti et al. (2023)). The fusion of masked autoencoder and contrastive learning is employed in both Voltron (Karamcheti et al. (2023)) and iBOT (Zhou et al. (2021)). The loss function achieves a balanced trade-off between these two aspects. To compare different pre-trained visual representations, benchmarks are established by CORTEXBENCH (Majumdar et al. (2023)) and EmbCLIP (Khandelwal et al. (2022)) to assess which model could provide a better ”artificial visual cortex” for manipulation tasks. However, the models included in these benchmarks are still not comprehensive enough. The aforementioned pre-trained visual representations mainly involve the extraction of features from 2D images. The experience of learning representations on 2D images can also be extended to other modalities. For the object point cloud modality, ULIP (Xue et al. (2023a)) and ULIP2 (Xue et al. (2023b)) employ contrastive learning to align features between point clouds and text-images. Point-BERT (Yu et al. (2022)) uses the masked autoencoding method to learn point cloud features by reconstructing point clouds. In the haptic modality, MOSAIC (Tatiya et al. (2023)) utilizes contrastive learning to train the haptic encoder.
As for segmentation, SAM (Kirillov et al. (2023)) develops a transformer-based architecture and creates the largest segmentation dataset, with over 1 billion masks from 11 million images. The model is adaptable and enables zero-shot transfer to new tasks and image distributions. Fast-SAM (Zhao et al. (2023b)) and Faster-SAM (Zhang et al. (2023a)) aim to improve the training and inference speed of the network by enhancing its network structure. TAM (Yang et al. (2023a)) merges SAM (Yang et al. (2023a)) and XMem (Cheng and Schwing (2022)) for high-performance interactive tracking and segmentation in videos.
As for detection, traditional detection models are usually confined to a narrow range of semantic categories because of the cost and time involved in gathering localized training data within extensive or open-label domains. However, advancements in language encoders and contrastive image-text training enable open-set detection. Researchers integrate language into a closed-set detector to generalize open-set concepts, detecting various classes through language generalization despite being trained solely on existing bounding box annotations, such as OWL-ViT (Minderer et al. (2022)), Grounding-DINO (Liu et al. (2023e)), OVD (Zareian et al. (2021)), ViLD (Gu et al. (2021)), DetCLIP (Yao et al. (2022a)).
The method for open-set detection on 2D images can be extended to the research direction of open-set detection on 3D point clouds. PointCLIP (Zhang et al. (2022)) utilizes pre-trained CLIP to extract multi-view depth image features of point cloud, then compares the extracted features with textual features to identify the point cloud category.
Summary
@@ 4.2 Object Recognition
Following Fig. 4. As for object affordance, current research explores LLMs’ zero/few-shot capability in providing object part-level prior knowledge via text, aiding in generating affordance maps or grasp poses. Direct use of foundation models for affordance is possible but still shows instability, with unclear performance boundaries. As for object recognition, representation learning facilitates the learning of the multimodal encoder, aligning features from various modalities to the text domain, aiding the model’s cognition, akin to how humans learn and think through words. The learning of the encoder promotes open-set perception tasks, such as open-set detection, open-set segmentation.
@@
Figure 4: Foundation Models for Pre-conditions Detection. As for object affordance, the main approaches of task-oriented grasp are supervised learning and reinforcement learning. Both methods utilize LLM to generate object part-level description and desired affordance description in task instruction, then fuse tokens and features into the original network through language encoder and image encoder to output task-oriented grasp pose (Tang et al. (2023); Ren et al. (2023b)). In reinforcement learning, it is possible to choose between a LLM language encoder with a custom-designed image encoder, or a VLM language encoder with a VLM image encoder. When selecting the LLM language encoder with a custom image encoder, the LLM language encoder should be frozen, and the custom image encoder should be trained (Ren et al. (2023b)). When using the VLM language encoder with the VLM image encoder, both encoders are typically frozen (Xu et al. (2023a)). Direct using foundation method utilizes LLM to generate object part-level description and desired affordance description according to task instruction. VLM marks out the part of the object to grasp in the image based on the description (Liu et al. (2023c)). As for object recognition, the representation learning methods in state perception mainly include contrastive learning (Radford et al. (2021)), distillation-based learning (Caron et al. (2021)), and masked autoencoder learning (Radosavovic et al. (2023)). Masked autoencoding methods prioritize low-level spatial aspects, sacrificing high-level semantics, whereas contrastive learning methods focus on the inverse, the fusion of masked autoencoder and contrastive learning is employed in both Voltron (Karamcheti et al. (2023)) and iBOT (Zhou et al. (2021)). Multimodal representation learning focuses primarily on multimodal alignment (Xue et al. (2023b); Tatiya et al. (2023)). Training the encoder with large-scale data and parameters has facilitated open-set perception, including tasks such as open-set detection, open-set segmentation. For instance, SAM (Kirillov et al. (2023)) utilizes the MAE (He et al. (2022)), ViLD (Gu et al. (2021)) employs the CLIP (Radford et al. (2021)).
@@ 5 Hierarchy of skills
In the last two decades, research on skill hierarchy has led to powerful domain-independent planners and various real-world applications (Silver et al. (2022)). Models for skill hierarchy can be trained using text or videos, similar to how humans learn assembly procedures from instructional manuals or tutorial videos. As for tutorial videos, VLaMP (Patel et al. (2023) and SeeDo (Wang et al. (2024a)) train models to understand human video operations and HourVideo (Chandrasegaran et al. (2024)) proposes a benchmark dataset specifically designed for hour-long video-language understanding. Currently, the text-based approach dominates the skill hierarchy domain. PDDL, a Lisp-like language (Silver et al. (2022)), is commonly used to address skill hierarchy issues. However, as LLMs excel in natural language tasks and PDDL is not a natural language, researchers are exploring how LLMs can be employed by robot for skill hierarchy tasks (Vemprala et al. (2023); Jansen (2020); Driess et al. (2023)). Additionally, various benchmarks such as PlanBench (Valmeekam et al. (2023)) and CALVIN (Mees et al. (2022)) can assess the planning and reasoning capability of LLMs.
@@
Regarding user preferences, TidyBot (Wu et al. (2023b)) trains LLM by collecting users’ preference data, enabling the trained LLM to choose behaviors that better align with user preferences. As for safety regulations, Yang et al. (2023d) incorporate ISO 61508, a global standard for safely deploying robots in industrial factory settings, into the constraints of the action sequence generation. As for the skill library, BOSS (Zhang et al. (2023c)) suggests using LLMs’ rich knowledge to guide skills chaining in the skill library, aiming to create new skills through combinations. RoboGen (Wang et al. (2023d)) employs generative models to create new skill task scenarios, then utilizes either reinforcement learning or gradient optimization methods to automatically learn new skills based on the reward function generated by the LLM. As for action execution feedback, REACT (Yao et al. (2022b)), COWP (Ding et al. (2022)), LLM-Planner (Song et al. (2023)), CoPAL (Joublin et al. (2023)) provide feedback on robot action execution to LLMs. This allows LLMs to adjust action sequences based on execution status, creating a closed-loop process for generating action sequences.
As for the limitation of scene grounding, LLMs need to inquire about the scene representation to determine the availability, relationship and location of objects. NLMap (Chen et al. (2023a)) proposes an open-vocabulary, queryable semantic representation map built on ViLD and CLIP. This map outputs the pose of related objects based on task instruction, which are then handed over to the LLM for planning. Text2Motion (Lin et al. (2023)) incorporates a geometric value function on top of the value function, enabling the robot to select actions that adhere to geometric constraints based on scene descriptions. Xu et al. (2023b) explore the possibility of teaching robots to creatively utilize tools within scenarios, which involve implicit physical limitations and require long-term planning. VILA (Hu et al. (2023a)) seamlessly incorporates perceptual data into ChatGPT-4V for its reasoning and planning processes, facilitating a deep comprehension of common sense knowledge within the visual domain, encompassing spatial arrangements and object characteristics. PHYSOBJECTS (Gao et al. (2023)) fine-tunes a VLM to enhance its understanding of physical object attributes, such as material. This integration of a physically informed VLM into an interactive framework with a LLM enhances task planning performance in tasks incorporating instruction related to physical object attributes. SpatialVLM (Chen et al. (2024b)) and 3D-LLM (Hong et al. (2023a)) utilizes a 2D pre-trained VLM to train on collected 3D datasets, enhancing capabilities related to 3D tasks while maintaining the abilities of previous tasks. Some methods, based on 3D scene graphs, provide 3D scene information for high-level planning, such as Zhang et al. (2021); Feng et al. (2023).
@@
6 State
The State module focuses on perceiving the environment, objects, and robot states. Section. 4 introduces low-level perception methods. This section explains high-level approaches for pose estimation and scene reconstruction.
Deep learning-based object pose estimation methods can be categorized into instance-level, category-level, and unseen object methods based on the problem formulation (Liu et al. (2024c)). Instance-level pose estimation methods are designed to estimate the pose of specific object instances. These methods work with a limited number of objects that are fully available in both the training and testing sets. Category-level methods enable the model to estimate poses for novel object instances within the same category. However, obtaining category-level training data is challenging due to the need for additional pose canonicalization and examination steps (Wen et al. (2024a)). However, these methods still struggle with generalizing to unseen object categories. In response, recent approaches have been proposed to address unseen object pose estimation (Liu et al. (2024c)).
There are some research integrating foundation models into pose estimation. As for category-level, OV9D (Cai et al. (2024)) proposes a framework for open-set category-level 9D object pose and size estimation. The key idea is to use DINO to extract features from masked RGB images. Then, it utilizes the Stable Diffusion (Rombach et al. (2022)) to generate a normalized object coordinate space (NOCS) map. The post-fitting algorithm uses the normalized object coordinate space map and the scene depth map to estimate the object’s size and pose. For unseen object pose estimation methods, current approaches mainly rely on open-set segmentation and detection capabilities to apply a mask or bounding box to unseen objects. This prepares the data for subsequent pose estimation algorithms (Lin et al. (2024b); Ausserlechner et al. (2024); Wen et al. (2024a)).